# Double Spending Analysis of Nakamoto Consensus for Time-Varying Mining Rates with Ruin Theory

### v1

@@ I Introduction
Recently, there has been a surge of interest in the investigation of the security-latency trade-off in PoW blockchains in the presence of network delays[12, 13, 14, 15, 16, 17]. In the so-called bounded delay model, all honest participants are guaranteed to receive a mined block after a bounded delay $\Delta$, for which the security-latency problem has been studied extensively in [18, 19, 20, 14] for continuous slots. These models are concerned with giving theoretical guarantees, therefore, they consider the case where the network delay experienced by most honest blocks is exactly $\Delta$, which we call fixed-delay model in this paper. In reality, however, some participants will receive a mined block soon after the mining instance and they can start to mine on a legitimate block right away, whereas a fraction of the participants will receive the mined block relatively late, creating forks. Out of an experimental study of information propagation in the Bitcoin network, [2] obtains an empirical probability density function of time until a node receives a block with a median of 6.5 seconds, mean of 12.6 seconds, and $5\%$ of the nodes receiving the block after 40 seconds. Further, [21] monitors the peer-to-peer network of Bitcoin and reports mean delay between the first observance of an inventory message and its relay by each particular peer the server is connected to. Both [2] and the raw data publicly available at [21] suggest that simply assuming zero-delay or fixed-delay may not reflect the reality. Zero-delay model neglects the situations where the network is congested and delays are high in general. Fixed-delay model, on the other hand, provides dramatically pessimistic results.
Another issue with fixed-delay results in [12, 18, 22, 19, 20, 14] is that they ignore the effects of the adversarial strategies on the block interval adjustment rule. It is shown in [23, Appendix A] that adversary should not publish any block before the attack. This makes sure that the block production rate has to be sustained by honest nodes, which in turn implies that the block puzzles (i.e., target thresholds) are easier than if adversary also contributed to the main chain. As a result, after the attack starts, the number of blocks mined during a delay interval, i.e., fork rate, increases, resulting in a higher double spending probability. Although the adversary considered in the works mentioned above does not contribute to the main chain before the attack, the immediate implication of this, i.e., an easier target threshold, is neglected. There has also been a recent interest to analyze the Nakamoto consensus under random network delays [24, 25, 26, 27]. Although the delay period is assumed to be exponentially distributed in [25], the authors also pessimistically assume that no node receives a block during the delay, and the results are then transferred to queuing theory and connected to the throughput. [26] generalizes the pessimistic models above to any random delay distribution and connects the results to the gap game [28, 29].
@@
Motivated by the work in [2], we attempt to model the honest mining rate during block delivery periods by an arbitrary function of time, which can be obtained from experimental data, to more accurately obtain the double spend probabilities. To the best of our knowledge, there is no work to model the honest mining rate as a function of time, as we do in this paper, for more accurate estimation of the double spend probabilities under variable network delays as well as dynamic participation of honest miners such as the gap game situations. However, the attacker’s mining rate is assumed to be fixed in this paper and the balanced attacks that can effectively increase it during block delivery periods are left for future research.
@@
  • The ruin-theoretical model we propose in this paper requires the first $k$ probability masses as well as the mean, of the distribution of blocks mined by the adversary during an honest inter-mining period, which is denoted by the non-negative integer-valued random variable $\Phi$. We propose to fit a matrix exponential distribution to the honest inter-mining times when the honest mining rate is time-varying, which subsequently allows us to obtain the probability mass function (pmf) and the mean of $\Phi$, employing a matrix-analytical algorithm, that is computationally stable and efficient.
  • The impact of private mining on block interval adjustment has been ignored in [18, 22, 19, 20, 25, 26]. As the number of blocks that can be mined during a delay interval, called the fork rate, has an impact on double spending, the mining rate increases when the adversary does not publish any blocks, e.g., private attack. Further, dynamic participation of honest miners also effects the block intervals and the double-spending probabilities. We consider this effect for the first time in our model.
@@ II-A ME-fication and Approximating Deterministic Variables
Phase-type (PH-type) distributions are very commonly used for modeling i.i.d. non-exponential inter-arrival and/or service times in queuing systems [32, 33, 34]. To describe a PH-type distribution, a Markov process is defined on the state-space $\mathcal{S}=\{1,2,\ldots,K,K+1\}$ with one absorbing state $K+1$, initial probability vector $v$ of size $1\times K$, and an infinitesimal generator of the form,
@@
where the sub-generator $T$ is $K\times K$, and $h$ is a column vector of size $K$ such that $h=-Te$, where $e$ denotes a column vector of ones of appropriate size. Note that the diagonal elements of the sub-generator $T$ are strictly negative, and its off-diagonal elements are non-negative. The distribution of time till absorption into the absorbing state $K+1$, is called PH-type characterized with the pair $(v,T)$, i.e., $X\sim PH(v,T)$ with order $K$. The cumulative distribution function (cdf) and the probability density function (pdf) of $X\sim PH(v,T)$ can be written as,
@@
which is a strictly proper rational function of $s$ with degree $K$ and $h=-Te$. Despite the lack of stochastic interpretations of the row vector $v$ and the matrix $T$, ME distributions can also be used in place of PH-type distributions using the same computational techniques and algorithms [38, 39, 40, 41].
A deterministic variable can be approximated by a PH-type, or more generally by an ME distribution. If $X$ is deterministic with $X=\delta$, its cdf is a unit step function located at $\delta$. In the Erlangization method, see for example [42, 43], $X=\delta$ is approximated by an Erlang-$K$ distribution with order $K$, i.e., $X^{(K)}\sim PH(v^{(K)},T^{(K)})$, denoted by $\text{ERL}[K,\delta]$, where
$v^{(K)}=$ $\begin{pmatrix}1&0&0&\cdots&0\end{pmatrix},$ (4)
@@
In the alternative ME-fication method, a family of ME distributions, called concentrated ME distributions (CME), are proposed in [44] for approximating deterministic quantities, and the parameters of the CME providing the minimal scv values are obtained numerically. The main result of [44] is that the asymptotic behavior of the scv of the proposed CME distribution of order $K$ is $\frac{2}{K^{2}}$. Fig. 1 depicts the cdf obtained with the approximate CME distribution, denoted by $\text{CME}[K,\delta]$ proposed in [44] for odd values of $K$ with rational mgfs when $\delta=2$ and $K\in\{7,17,27,37,47\}$. The accuracy of this approximation improves quadratically as $K$ increases, for example, when $K=47$, the scv drops down to 0.000882. With the order of the approximating distribution being fixed, ME-fication provides better fits than Erlangization [45].
@@ II-C Lindley Equation
$p_{Q}(i)$ $=\lim_{i\rightarrow\infty}\mathbb{P}(Q_{i}=n)=\frac{\sum_{j=0}^{i-1}p_{Q}(j)\bar{F}_{\Phi}(i-j)}{p_{\Phi}(0)},\ i\geq 1.$ (13)
We also let ${F}_{Q}$ and $\bar{F}_{Q}$ denote the cdf and ccdf of the random variable $Q$. The $k$-partial pgf of $Q$ is,
$G^{(k)}_{Q}(z)$ $=\sum_{n=0}^{k-1}p_{Q}(n)z^{n}.$ (14)
@@ II-D Discrete-Time Risk Problem
A specific discrete-time insurance risk problem studied in [31] considers the following discrete-time surplus process $U_{n}$, $n=1,2,\ldots$,
@@
where $u=U_{0}$ is the insurer’s initial surplus, $\Phi_{i}$ denotes the insurer’s aggregate claim amount in interval $i$, $\Phi_{i}$ are i.i.d., i.e., $\Phi_{i}\sim\Phi$, taking non-negative integer values. The insurer’s premium income in one interval is one. For this surplus process, ultimate ruin occurs if the surplus ever falls to zero or below. We let the random variable,
@@ III System Model
As the attacker can mine on its own private chain as long as it is longer than the honest chain, eventually it can build a private branch long enough to undo $tx$ with a $tx^{\prime}$ even before the arrival of transaction $tx$. Therefore, we consider the pre-mining attacks studied in [49] where the arrival of $tx$ is random and the adversary mines a private block whenever the private branch is longer than the honest branch. When the honest chain is longer, then the attacker starts a new branch on the honest chain until the arrival of $tx$. Therefore, there is a possibility that the adversary chain may be longer than the honest chain when the transaction $tx$ appears in the mempool, which we call the lead in this paper, and find its distribution.
Let $\Theta_{i}$ denote the inter-mining times of honest blocks $i$ and $i+1$; see Fig. 2. We assume that $\Theta_{i}$s are i.i.d. and $\Theta_{i}\sim\Theta$ with pdf $f_{\Theta}(x)$, $x\geq 0$, and corresponding mgf $M_{\Theta}(s)=\mathbb{E}[\mathrm{e}^{s\Theta}]$. Also, let $\Phi_{i}$ denote the number of blocks mined by the adversary during the inter-mining time $\Theta_{i}$. Since the mining rate of the adversary is fixed in this paper, $\Phi_{i}\sim\Phi$ with pmf $p_{\Phi}(n)$, $n=0,1,\ldots$ and corresponding pgf $G_{\Phi}(z)=\mathbb{E}[z^{\Phi}]$. Clearly, the network delay $\Delta$ and the hashrate function $\alpha(t)$, $0\leq t\leq\Delta$, have impact on the distribution of $\Phi$ which will be explored later in the paper. The sub-case $\Delta=0$ is the zero-delay model investigated in [1, 11] whereas the sub-case of $\alpha(t)=0$, $0\leq t\leq\Delta$, is the fixed-delay model, in which all miners receive the mined block after a fixed delay $\Delta$, and therefore, when the mined block is being announced, the hashrate $\alpha$ is assumed to be non-usable for mining a new block, instead will be wasted towards generation of forks; see for example [19, 22]. In this paper, we take a more general approach by providing a novel ruin-theoretical model for double spending when the hashrate function is an arbitrary function of time, covering the zero-delay and fixed-delay models as sub-cases.
We assume $\mathbb{E}[\Phi]<1$, a $\mathcal{T}$-block interval rule and the $k$-deep confirmation rule. The condition $\mathbb{E}[\Phi]<1$ makes sure that the double spending probability is not trivially $1$. $\mathcal{T}$-block interval rule adjusts the puzzle difficulty such that the average time for mining a block remains $\mathcal{T}$ seconds on average. According to the $k$-deep confirmation rule, a block (along with all the transactions inside it) is said to be confirmed if it is part of the longest chain and at least $k-1$ blocks have been mined on top of it. The safety violation, or double spending, refers to the event that a transaction $tx$ is confirmed according to the $k$-deep confirmation rule in the view of an honest miner and later gets replaced (or discarded) with another conflicting transaction $tx^{\prime}$ in the same honest view or another honest view. In this paper, we provide a matrix-analytical method, which requires computationally efficient and stable matrix-vector operations only, for obtaining the double spending probability under a certain private attack.
@@ IV The Analytical Method
In the zero-delay model, the propagation delay $\Delta=0$. Thus, once block $i$ is mined, the full hashrate $\alpha$ is used to mine block $i+1$. Hence, block inter-arrival times are exponentially distributed with parameter $\alpha$, i.e., mean $\alpha^{-1}$, giving rise to the following rational mgf,
@@
In the fixed-delay model, the propagation delay is fixed as $\Delta$, and all miners receive the mined block after the fixed delay $\Delta$. Thus, when the mined block is being announced, the honest mining power is assumed to be non-usable for mining a new block, giving rise to the following mgf $M_{\Theta}(s)$,
$M_{\Theta}(s)$ $=\mathrm{e}^{\Delta s}\ \frac{\alpha}{\alpha-s}.$ (23)
The expression above is not rational. Thus, we propose to use the $\text{CME}[K,\Delta]$ ME-fication method which uses a CME $X\sim ME(v^{(1)},T^{(1)})$ with mean $\Delta$ and order $K$ so that $M_{\Theta}(s)$ is written in the general form (3) with,
@@
For the exponential delay model of [25], the network delay is assumed to be between $[0,\Delta_{h}]$ for blocks at height $h$ with $\Delta_{h}$, $h\geq 1$ being i.i.d. exponential random variables as opposed to being deterministic. For the case of $\Delta_{h}$ having a general ME distribution characterized with the pair $\Delta_{h}\sim ME(v^{(1)},T^{(1)})$, then $M_{\Theta}(s)$ is written in the form (3) where the parameters of this expression, namely $v,T$ and $h$ are given as in (24), (27) and (28), respectively.
@@
As we assume a private attack with pre-mining, the adversarial blocks do not contribute to the block intervals. Thus, $\mathcal{T}$-block interval rule implies that $\mathbb{E}[\Theta]=\mathcal{T}$ has to be satisfied. Note that,
@@
A simple method to achieve $-vT^{-1}e=\mathcal{T}$ is to first assume $\alpha^{(1)}=1/\mathcal{T}$, find associated $T^{(1)}$ and $\mathbb{E}_{\Theta}^{(1)}$, then repeat with $\alpha^{(i+1)}=\alpha^{(i)}\times\mathbb{E}_{\Theta}^{(i)}/\mathcal{T}$ accordingly until a predetermined accuracy (we pick $10^{-4}$) is reached. It is also possible to use bisection search on $\alpha$. Throughout the paper, we assume that $M_{\Theta}(s)$ is given as in (37) for describing the inter-mining times on the honest chain, which covers all the scenarios described in this section, namely zero-delay, fixed-delay, random-delay and variable-delay models.
In the second step, we obtain the $k$-partial pgf and the mean of the random variable $\Phi$ which stands for the number of blocks mined by the adversary in an inter-mining interval. The pgf of $\Phi$, denoted by $G_{\Phi}(z)$, can be written from (10),
$G_{\Phi}(z)$ $=M_{\Theta}(s)\Big{|}_{s=\beta(z-1)}$ (39)
$=\left(-v(sI+T)^{-1}h\right)\bigg{|}_{s=\beta(z-1)}$ (40)
@@
with the square matrix $A$ being of size $m$. From (9), the pmf $p_{\Phi}(n)$ can then be written as in (7). Thus, the first $k$ probability masses, namely $p_{\Phi}(n)$, $10\leq n\leq k-$ 1, can be obtained by the following,
@@
To calculate the double spending probability, we will work with its complementary event, hence, we are not interested in the rest of the probability masses, namely $p_{\Phi}(n),n\geq k$. The mean of $\Phi$, $\mathbb{E}[\Phi]$, can be written as,
@@
The double spending probability is described in line with the diagram given in Fig. 2. We assume that the particular transaction $tx$ is included in honest block at height $r+1$, where $r\gg 1$, which is mined at time $\tau_{1}$, where $\tau_{0}$ denotes the epoch where the honest block at height $r$ is mined. We also let $\tau_{i}+\Delta$, $i=1,2,\ldots$ denote the time when the block containing $tx$ is $i$-deep in all honest views. Once the transaction $tx$ appears at time $t$, where $\tau_{0}<t<\tau_{1}$, the adversary starts to mine only on its own chain, called the private chain. According to the $k$-deep confirmation rule, at time $\tau_{k}+\Delta$, the transaction $tx$ is said to be confirmed in all honest views.
@@
Let $t^{-}$ denote the time just before the instant $t$. Using [26, Lemmas 1-2], we have,
@@
since $A_{t}$ is monotone increasing and $H_{t}^{v}$ has jumps at time instances ${(\tau_{k+i}+\Delta)}$ for some honest view $v^{*}$ which is yet to observe the honest block at height $r+k+i$. The honest view $v^{*}$ has the same function as the weakest possible view explained in [20, Section III.A]. Thus, we can focus on safety violations at the embedded instants $(\tau_{k+i}+\Delta)^{-}$, $i\geq 1$, as opposed to arbitrary time points. For this purpose, we let the random variable $\kappa$ denote the first such epoch
@@
A safety violation at an arbitrary time $t$ immediately implies a safety violation at the next embedded instant. Therefore, we conclude that the quantities $p$ and $q$ are the same. Thus, in the rest of this paper, we will focus our attention to obtaining the probability $q$ in (53).
Similar to [18, 22, 19, 25, 20, 26], we divide our analysis into three phases: pre-mining phase, confirmation interval and post-confirmation race. In the pre-mining phase, the adversary tries to build a private chain to gain some advantage (lead) over the honest chain. When the lead is less than or equal to zero, the adversary will always build a new private block on top of the honest chain in the pre-mining phase, otherwise, it will continue mining on its own private chain. Once the pre-mining phase is over, the adversary will always build on its own private chain. Here, we note a subtle technical difference in our analysis from the works of [18, 22, 19, 25, 20, 26]. In those works, the authors assume to observe the lead at a particular instant $t$ when $tx$ arrives at the mempool, and then start the race of mining again. As block arrivals are modeled exponential, inspection paradox gives rise to an over-estimation of the double-spending probability in those papers as they consider worst case scenarios and want to give theoretical upper bounds. Moreover, one could note that the overestimation is not necessary as honest miners cannot observe the private chain which would lead to the inspection paradox. As our goal here is to provide results with more realistic settings, we consider the renewal interval containing $t$ is the same as any other interval. In essence, the difference of models involving the inspection paradox correspond to an additional adversarial advantage of $G$ in the distribution of the adversarial lead, where $G$ is geometrically distributed with success probability $\frac{\alpha}{\alpha+\beta}$.

### v2

@@ I Introduction
Recently, there has been a surge of interest in the investigation of the security-latency trade-off in the presence of network delays[12, 13, 14, 15, 16]. In the so-called bounded delay model, all honest participants are guaranteed to receive a mined block after a bounded delay $\Delta$, for which the security-latency problem has been studied extensively in [17, 18, 19, 20]. These models are concerned with giving theoretical guarantees, therefore, they consider the case where the network delay experienced by most honest blocks is exactly $\Delta$, which we call fixed-delay model in this paper. In reality, however, some participants will receive a mined block soon after the mining instance and they can start to mine on a legitimate block right away, whereas a fraction of the participants will receive the mined block relatively late, creating forks. Out of an experimental study of information propagation in the Bitcoin network, [2] obtains an empirical probability density function of time until a node receives a block with a median of 6.5 seconds, mean of 12.6 seconds, and $5\%$ of the nodes receiving the block after 40 seconds. Both [2] and the raw data of the inventory messages publicly available at [21] suggest that simply assuming zero-delay or fixed-delay may not reflect the reality.
There has also been a recent interest to analyze the Nakamoto consensus under random network delays [22, 23, 24, 25]. The delay period is assumed to be exponentially distributed in [23] and the results are then transferred to queuing theory and connected to the throughput. [24] generalizes the pessimistic models above to any random delay distribution and connects the results to the mining gap [26, 27].
Another issue with fixed-delay results in [12, 17, 28, 18, 19, 20] is that they ignore the effects of the adversarial strategies on the block interval adjustment rule. It is shown in [29, Appendix A] that the adversary should not publish any block before the attack. For example, if the adversary mines on a different cryptosystem temporarily, the puzzle difficulty decreases similar to the coin-hopping attacks [30, 31]. Then, when the adversary mounts the attack, during a delay interval, the number of adversarial blocks mined on new heights increases while honest miners are wasting their hashpower towards forks. As a result, forkrate and double spending probability increase.
@@
Motivated by the work in [2], we attempt to model the honest mining rate during block delivery periods by an arbitrary function of time, which can be obtained from experimental data, to more accurately obtain the double spend probabilities. To the best of our knowledge, there is no work to model the honest mining rate as a function of time, as we do in this paper, for more accurate estimation of the double spend probabilities under variable network delays. However, the attacker’s mining rate is assumed to be fixed in this paper and the balanced attacks that can effectively increase it during block delivery periods are left for future research.
@@
  • The ruin-theoretical model we propose in this paper requires the first $k$ probability masses as well as the mean, of the distribution of blocks mined by the adversary during an honest inter-mining period, which is denoted by the non-negative integer-valued random variable $\Phi$. We propose to fit a matrix exponential distribution to the honest inter-mining times when the honest mining rate is time-varying, which subsequently allows us to obtain the probability mass function (pmf) and the mean of $\Phi$, employing a matrix-analytical algorithm, that is computationally stable and efficient. The impact of block interval adjustment rule has been ignored in [17, 28, 18, 19, 23, 24]. When the adversary does not publish any blocks temporarily, the block puzzles become easier and the number of blocks that can be mined during a delay interval increases which has an impact on double spending. Further, dynamic participation of honest miners also effects the block intervals and the double spending probabilities. We consider this effect for the first time in our model.
@@ II-A ME-fication and Approximating Deterministic Variables
Phase-type (PH-type) distributions are very commonly used for modeling in queuing systems [34, 35, 36]. A Markov process is defined on the state-space $\mathcal{S}=\{1,2,\ldots,K,K+1\}$ with one absorbing state $K+1$, initial probability vector $v$ of size $1\times K$, and an infinitesimal generator of the form,
@@
where the sub-generator $T$ is $K\times K$, and $h$ is a column vector such that $h=-Te$, where $e$ is a column vector of ones. The diagonal elements of are strictly negative, and its off-diagonal elements are non-negative. The distribution of time till absorption into state $K+1$, is called PH-type characterized with the pair $(v,T)$, i.e., $X\sim PH(v,T)$ with order $K$. The cumulative distribution function (cdf) and the probability density function (pdf) of $X\sim PH(v,T)$ can be written as,
@@
which is a strictly proper rational function of $s$ with degree $K$ and $h=-Te$. ME distributions can also be used in place of PH-type using the same computational techniques [40, 41, 42, 43].
A deterministic variable, $X$, can be approximated by a ME distribution. In the Erlangization method [44, 45], $X=\delta$ is approximated by an Erlang-$K$ distribution with order $K$, i.e., $X^{(K)}\sim PH(v^{(K)},T^{(K)})$, where $v^{(K)}=\begin{pmatrix}1&0&0&\cdots&0\end{pmatrix}$ and
@@
In the alternative ME-fication method, concentrated ME distributions (CME) are proposed in [46] for approximating deterministic quantities. The asymptotic behavior of the scv of the proposed CME distribution of order $K$ is $\frac{2}{K^{2}}$. Fig. 1 depicts the cdf obtained with the approximate CME distribution, denoted by $\text{CME}[K,\delta]$ proposed in [46] for odd values of $K$ with rational mgfs when $\delta=2$. The accuracy of this approximation improves quadratically as $K$ increases. With the order of the approximating distribution being fixed, ME-fication provides better fits than Erlangization [47].
@@ II-C Lindley Equation
$p_{Q}(n)$ $=\lim_{i\rightarrow\infty}\mathbb{P}(Q_{i}=n)=\frac{\sum_{j=0}^{n-1}p_{Q}(j)\bar{F}_{\Phi}(n-j)}{p_{\Phi}(0)},\ n\geq 1.$ (11)
@@ II-D Discrete-Time Risk Problem
Recently, insurance risk problems were used in finding probabilities of successful withholding attacks and associated profits in the blockchain context [51, 52]. Here, we consider a specific discrete-time insurance risk problem studied in [33], i.e., the discrete-time surplus process $U_{n}$, $n=1,2,\ldots$,
@@
where $u=U_{0}$ is the insurer’s initial surplus, $\Phi_{i}$ denotes the insurer’s aggregate claim amount in interval $i$, $\Phi_{i}\sim\Phi$ are i.i.d. taking non-negative integer values. The insurer’s premium income in one interval is one. Ultimate ruin occurs if the surplus ever falls to zero or below. Let the random variable,
@@ III System Model
We consider the pre-mining attacks [53] where the arrival of $tx$ is random and the adversary mines a private block whenever the private branch is longer than the honest branch. When the honest chain is longer, then the attacker starts a new branch on the honest chain until the arrival of $tx$. Therefore, there is a possibility that the adversary chain may be longer than the honest chain when the transaction $tx$ appears in the mempool, which we call the lead in this paper, and find its distribution.
Let $\Theta_{i}$ denote the inter-mining times of honest blocks $i$ and $i+1$; see Fig. 2. We assume that $\Theta_{i}$s are i.i.d. and $\Theta_{i}\sim\Theta$ with pdf $f_{\Theta}(x)$, $x\geq 0$, and corresponding mgf $M_{\Theta}(s)=\mathbb{E}[\mathrm{e}^{s\Theta}]$. Also, let $\Phi_{i}$ denote the number of blocks mined by the adversary during the inter-mining time $\Theta_{i}$. Since the mining rate of the adversary is fixed in this paper, $\Phi_{i}\sim\Phi$ with pmf $p_{\Phi}(n)$, $n=0,1,\ldots$ and corresponding pgf $G_{\Phi}(z)=\mathbb{E}[z^{\Phi}]$. Clearly, the network delay $\Delta$ and the hashrate function $\alpha(t)$, $0\leq t\leq\Delta$, have impact on the distribution of $\Phi$.
We assume $\mathbb{E}[\Phi]<1$, a $\mathcal{T}$-block interval rule and the $k$-deep confirmation rule. The condition $\mathbb{E}[\Phi]<1$ makes sure that the double spending probability is not trivially $1$. $\mathcal{T}$-block interval rule adjusts the puzzle difficulty such that the average time for mining a block remains $\mathcal{T}$ seconds on average. According to the $k$-deep confirmation rule, a block (along with all the transactions inside it) is said to be confirmed if it is part of the longest chain and at least $k-1$ blocks have been mined on top of it. The safety violation, or double spending, refers to the event that a transaction $tx$ is confirmed according to the $k$-deep confirmation rule in the view of an honest miner and later gets replaced (or discarded) with another conflicting transaction $tx^{\prime}$ in the same honest view or another honest view.
@@ IV The Analytical Method
In the zero-delay model, $\Delta=0$, once block $i$ is mined, the full hashrate $\alpha$ is used to mine block $i+1$. Hence, block inter-arrival times are exponentially distributed with parameter $\alpha$, giving rise to the following rational mgf,
@@
In the fixed-delay model, the propagation delay is fixed as $\Delta$, and all miners receive the block after $\Delta$ time. During the delay, the honest mining power is assumed to be non-usable for mining a new block, giving rise to the following mgf $M_{\Theta}(s)$,
$M_{\Theta}(s)$ $=\mathrm{e}^{\Delta s}\ \frac{\alpha}{\alpha-s}$ (20)
which is not rational. Thus, we propose to use the $\text{CME}[K,\Delta]$ ME-fication method which uses a CME $X\sim ME(v^{(1)},T^{(1)})$ so that $M_{\Theta}(s)$ is written in the general form (3) with,
@@
For the exponential delay model of [23], the network delay is assumed to be between $[0,\Delta_{h}]$ for blocks at height $h$ with $\Delta_{h}$, $h\geq 1$ being i.i.d. exponential random variables. For the case of $\Delta_{h}$ having a general ME distribution characterized with the pair $\Delta_{h}\sim ME(v^{(1)},T^{(1)})$, then $M_{\Theta}(s)$ is written in the form (3) where the parameters of this expression, namely $v,T$ and $h$ are given as in (21), (24) and (25), respectively.
@@
If the adversary temporarily works on another system before it mounts the attack, $\mathbb{E}[\Theta]=\mathcal{T}$ needs to be satisfied where only $\alpha(t)$ determines the puzzle difficulties. Note that,
@@
A simple method to achieve $-vT^{-1}e=\mathcal{T}$ is to first assume $\alpha=1/\mathcal{T}$, find associated $T$ and $\mathbb{E}_{\Theta}$, then repeat with $\alpha\coloneqq\alpha\times\mathbb{E}_{\Theta}/\mathcal{T}$ accordingly until a predetermined accuracy (we pick $10^{-4}$) is reached. It is also possible to use bisection search on $\alpha$. We assume that $M_{\Theta}(s)$ is given as in (34), which covers all the scenarios described in this section, namely zero-delay, fixed-delay, random-delay and variable-delay models.
Next, the pgf of $\Phi$, $G_{\Phi}(z)$, can be written from (8),
$G_{\Phi}(z)$ $=M_{\Theta}(s)=\left(-v(sI+T)^{-1}h\right)\bigg{|}_{s=\beta(z-1)}$ (36)
@@
with the square matrix $A$ being of size $m$. From (7), the pmf $p_{\Phi}(n)$ can then be written as in (6). Thus, the first $k$ probability masses can be obtained by the following,
@@
We will work with the complementary event of double spending, hence, we are not interested in the rest of the probability masses $p_{\Phi}(n),n\geq k$. The mean of $\Phi$ can be written as,
@@
The double spending probability is described in line with the diagram given in Fig. 2. We assume that the particular transaction $tx$ is included in honest block at height $r+1$, $r\gg 1$, which is mined at time $\tau_{1}$, where $\tau_{0}$ denotes the epoch where the honest block at height $r$ is mined. We also let $\tau_{i}+\Delta$, $i=1,2,\ldots$ denote the time when the block containing $tx$ is $i$-deep in all honest views. Once the transaction $tx$ appears at time $t$, $\tau_{0}<t<\tau_{1}$, the adversary starts to mine only on its own private chain. According to the $k$-deep confirmation rule, at time $\tau_{k}+\Delta$, $tx$ is said to be confirmed in all honest views.
@@
Let $t^{-}$ denote the time just before the instant $t$. Using [24, Lemmas 1-2], we have,
@@
since $A_{t}$ is monotone increasing and $H_{t}^{v}$ has jumps at time instances ${(\tau_{k+i}+\Delta)}$ for some honest view $v^{*}$ which is yet to observe the honest block at height $r+k+i$. The honest view $v^{*}$ has the same function as the weakest possible view explained in [19, Section III.A]. Thus, we can focus on safety violations at the embedded instants $(\tau_{k+i}+\Delta)^{-}$, $i\geq 1$, as opposed to arbitrary time points. For this purpose, we let the random variable $\kappa$ denote the first such epoch
@@
A safety violation at an arbitrary time $t$ immediately implies a safety violation at the next embedded instant. Therefore, we conclude that the quantities $p$ and $q$ are the same. Thus, we will focus our attention to obtaining the probability $q$ in (49).
Similar to [17, 28, 18, 23, 19, 24], we divide our analysis into three phases: pre-mining phase, confirmation interval and post-confirmation race. In the pre-mining phase, the adversary tries to build a private chain to gain some advantage (lead) over the honest chain. When the lead is less than or equal to zero, the adversary will always build a new private block on top of the honest chain in the pre-mining phase, otherwise, it will continue mining on its own private chain. Once the pre-mining phase is over, the adversary will always build on its own private chain. Here, we note a subtle technical difference in our analysis from the existing works. There, the authors assume to observe the lead at a particular instant $t$ when $tx$ arrives at the mempool, and then start the race of mining again. As block arrivals are modeled exponential, inspection paradox gives rise to an over-estimation of the double spending probability in those papers as they consider worst case scenarios and want to give theoretical upper bounds. As our goal here is to provide results with more realistic settings, we consider the renewal interval containing $t$ is the same as any other interval.